OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 + 6x − 2. (1, 1) (−1, 1) (1, −5) (−1, −5)

OpenStudy (anonymous):

To find the vertex use the formula below to find the x vertex point of y = −3x^2 + 6x − 2 x=−b2a Plug our points in the the formula x=−(+62(−3) x=−6−6 x = 1 //This is our x vertex Now we have the x coord (1, ?) we need to find the y coord To find the y coord just plug our x coord, which is -1 into the equation y=−3x2+6x−2 y=−3(1)2+6(1)−2 y=−3+6−2 y=3−2 y =1 //This is our y vertex point Now we have x and y so our vertex is at (1, 1 Right or wrong???

OpenStudy (anonymous):

so the vertex formula is -b/2a, first of all

OpenStudy (anonymous):

so thats -6/-6, which equals 1

OpenStudy (anonymous):

yup you're right!!